• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 04:01:57 on November 18, 2008

    Entry number 249452

    keyword=target motion problem

    Around 11:00 pm when TO tried to move target from BeO to He3, it did not seem to move (from the camera) while the GUI shows it moved to a different position. TO called target-on-call (Yi) and Yi tried to move to another position. It appeared to not response correctly. I suggested to reboot ioc (similar situation happened before and reboot often solved problem). However, this time reboot did not work. We called tech-on-call, Heidi. She suggested we try to power cycle the controller. We (Aelxandre and I) went into the hall (in controlled access) did power cycle and reboot ioc a number of times, but did not fix the problem. Heidi came in. I went with Heidi in the hall. Heidi checked the controller, the relay box and all the connections. The controller has power but the ONLINE light not on. The IOC does talk to other devices, but not with the motion controller. Heidi replaced the motion controller with the one for left HRS angle camera, still not working. Heidi called in Jack and two of them have been working down there for about two hours. Jack was able to talk to the controller through with a laptop through a serial line, but ioc does not communicate with the controller. Jack think it probably is a software problem. With Jack's suggestion, I called MCC to get the software-on-call. David callled and discussed with Jack and did some check. The problem is still not fixed. Jack moved the target to the optics/BeO position and came out of the hall now. We are going back to beam permit. Will get tuning beam on the BeO to check position. And then do some low current test runs for the BB gas Cherenkov. Will wait to the morning to have Pam from the software group to work on the motion problem.



    A copy of this log entry has been emailed to: camsonne@jlab.org,gen@jlab.org,yqiang@jlab.org